# thanks: Corresponding author

### v1

@@ Abstract
This paper explores recommender systems in social networks which leverage information such as item rating, intra-item similarities, and trust graph. We demonstrate that item-rating information is more influential than other information types in a collaborative filtering approach. The trust graph-based approaches were found to be more robust to network adversarial attacks due to hard-to-manipulate trust structures. Intra-item information, although sub-optimal in isolation, enhances the consistency of predictions and lower-end performance when fused with other information forms. Additionally, the Weighted Average framework is introduced, enabling the construction of recommendation systems around any user-to-user similarity metric.
@@ I Introduction
Recommendation systems have become an essential component of online platforms, helping users navigate vast amounts of information and make decisions based on personalized suggestions. In social networks, these systems play a key role in suggesting products, content, friends, or pages to users based on their preferences, behaviors, and social connections. Social networks provide a rich source of information on user interactions and preferences, which can be leveraged to create more accurate and personalized recommendations. This type of recommendation system considers not only the past behaviors and preferences of users but also the behaviors and preferences of their connections or friends in the network.
@@
Trust is a crucial concept in social networks and plays a vital role in recommendation systems [1, 2, 3, 4, 5]. The abstraction of social connections as ’trust’ is essential because it allows the system to consider implicit behaviors and preferences of users and their connections. For example, a user liking or sharing content, or having many mutual connections, can indicate a level of trust or shared interest, even if the users are not explicitly connected or have not directly interacted. This abstraction helps in capturing the nuances of social interactions and enables the recommendation system to provide more personalized and accurate suggestions. Additionally, the concept of trust incorporates a user’s social circle’s opinions and preferences, making recommendations more robust and comprehensive [6]. In this paper, we investigate trust connections, and the implications they have on a user’s preferences and the robustness to adversarial attacks.
@@
There are several types of recommendation systems, including content-based, collaborative filtering, hybrid, and others. Content-based recommendation systems suggest items based on the characteristics of the items and the preferences of the users. On the other hand, collaborative filtering recommends items based on the preferences and behaviors of similar users [7]. Hybrid systems combine both approaches. Trust-based collaborative filtering is particularly interesting because it not only considers the preferences of similar users but also incorporates the trust relationships between users. This approach helps in mitigating some of the limitations of traditional collaborative filtering, such as the cold start problem and data sparsity, by leveraging the trust network to generate recommendations even when there is limited interaction data available [8]. This is a technique we will explore in this paper, but with a first principles approach to understand how to utilize these different methods and information forms best.
@@
Outline. The paper is organized in the following manner. Section II gives an overview of prior work, and section III outlines our contribution. After that, in section IV we introduce the methodology and in section V we discuss the single information-based algorithms and subsequently multiple information-based algorithms in section VI. Validation of the findings on more datasets is given in section VII and the impact of adversarial attacks is explored in section VIII. A discussion of our main findings is given in section IX.
@@ II.1 Traditional Approaches
Traditional trust-based recommendation systems mainly focus on incorporating explicit and implicit trust relationships into the recommendation process. For example, some approaches consider the ratings or preferences of a user’s trusted friends when generating recommendations. [9] delved into a trust-aware collaborative filtering approach. Their method distinctively pivots on explicit trust relationships. This means that the system directly harnesses expressed trust values (like friendship connections) to modulate the influence of a user’s peers on their recommendations.
@@ II.2 Cold-Start Approaches
Addressing the “cold start” problem remains a pivotal challenge for trust-based recommendation systems. The “cold start” issue surfaces when a new user joins the platform or a new item is added, and there’s minimal to no interaction data available for meaningful recommendation generation. In such scenarios, traditional collaborative filtering methods often fall short, as they rely predominantly on historical interaction data [12]. Various solutions have been proposed, including adaptive clustering and incremental learning [13], and a demographic approach based on finding similarities between old and new users [14].
Another porpular approach is leveraging the trust network [15]. The assumption here is that even if a user hasn’t interacted with many items, their trusted connections might have. By leveraging this network, systems can extrapolate the potential interests of the new user based on the preferences of their trusted peers. One particularly noteworthy approach to leverage the trust network against the “cold start” problem was conceived by [16]. They introduced TrustWalker, a random walk model grounded in the trust network. It commences its journey from the target user and propagates through their trusted connections, and with some probability, swapping to other similar items that may have more information. In this way, TrustWalker harnesses both explicit and implicit trust connections along with skipping to different items, which lets “cold start” users with little information still receive “acceptable” recommendations.
@@ II.3 Machine Learning Approaches
Machine learning, over the past decade, has revolutionized numerous domains with its ability to extract patterns from vast data sets. Traditional methods of collaborative filtering were primarily linear. However, the dynamics of user preferences, trust relationships, and item attributes often exhibit nonlinear patterns that require more sophisticated models for accurate representation [17]. Hence, machine learning is a widely used tool in recommendation systems [18]. Bayesian [19, 20] and decision tree [21, 22] approaches are mostly used due to their simplicity [23]. Moreover, gradient-decent-based [24], matrix factorization-based [25], and hidden Markov model-based [26] algorithms are also in use.
Neural collaborative filtering models underscore the potential of deep neural networks in handling the collaborative filtering challenge on implicit feedback data sets [27]. They propose three different ML-based models for developing recommendation systems and show the models compete with other state-of-the-art approaches. Moreover, these approaches have also been used to tackle the cold-start problem [28, 29].
@@ II.4 Adversarial Defense Approaches
Low-rank defenses [33] against adversarial attacks are one type of adversarial defense strategy used in recommendation systems. Low-rank defenses entail transforming the user-item matrix into a low-rank matrix to decrease the impact of adversarial attacks. This method has been demonstrated to be effective in fighting against several sorts of adversarial attacks, including poisoning attacks and model extraction attacks. In recommendation systems, generative adversarial networks (GANs) can detect and fight against adversarial attacks. GANs can be used to generate adversarial instances and teach the recommendation model to defend against them. This technique effectively defends against adversarial assaults such as poisoning and model extraction [31]. Another study by [34] focused on how to accurately portray the representative rating for an item to mitigate the risk of an adversary adding fake ratings. The paper found that placing weights on network-related factors leads to increased robustness to adversaries.
@@ V.1.3 Monte-Carlo Random Walk
We now move to investigate recommendation systems that consider the entire trust graph. The Monte-Carlo (MC) Random Walk model works by starting at the node we are trying to produce a rating for and then performing a random walk. As the distance from the original node, $d$, increases, so does the probability that a random rating is returned. This random walk continues until a node with a rating is reached, or this random rating is produced. We run this experiment $k$ times, for some fixed $k$, for each node and take the mean rating from the nodes reached via random walk.
@@ V.1.4 Jaccard Majority of Majorities (JMoM) and Weighted Average (Jaccard WA)
These recommenders work by considering every other node, which has rated item i, in the graph. We then perform a summation of the Jaccard values for every node corresponding to each rating value. In the Majority of Majorities version, we then take the rating that has received the highest contribution of Jaccard scores. However, in the Weighted Average version, we instead take the weighted average of these ratings, weighted by the contribution of Jaccard scores. We will call these techniques JMoM and WA moving forward. (These methods are inspired by the popular majority based opinion diffusion models, cf. [38, 39, 40].)
@@ VII Validation With Other Data Sets
Jaccard Item-Jaccard WA 1.00 0.26 0.66 0.08 0.53 0.28
@@
It appears the Jaccard Item-Jaccard WA recommendation algorithm performed best across all data sets with the lowest MAE value. The algorithm consistently outperformed other algorithms in each of the data sets - Epinions, FilmTrust, and CiaoDVD - with MAEs of 1.00, 0.66, and 0.53 respectively. This algorithm’s Mean MAE ($\mu_{MAE}$) was significantly lower than other algorithms.
@@
VIII Robustness Against Adversarial Attacks
@@ IX Discussion
Paras Stefanopoulos: Conceptualization, Formal Analysis, Methodology, Software, Validation, Writing - Original draft preparation. Ahad N. Zehmakan: Conceptualization, Methodology, Supervision, Validation, Writing - Review. Sourin Chatterjee: Software, Validation, Writing - Original draft preparation.

### v2

@@ Abstract
This paper explores recommender systems in social networks which leverage information such as item rating, intra-item similarities, and trust graph. We demonstrate that item-rating information is more influential than other information types in a collaborative filtering approach. The trust graph-based approaches were found to be more robust to network adversarial attacks due to hard-to-manipulate trust structures. Intra-item information, although sub-optimal in isolation, enhances the consistency of predictions and lower-end performance when fused with other information forms. Additionally, the Weighted Average framework is introduced, enabling the construction of recommendation systems around any user-to-user similarity metric. All the codes are publicly available on GitHub [50].
@@ I Introduction
Recommendation systems have become an essential component of online platforms, helping users navigate vast amounts of information and make decisions based on personalized suggestions. In social networks, these systems play a key role in suggesting products, content, friends, or pages to users based on their preferences, behaviors, and social connections. Social networks provide a rich source of information on user interactions and preferences, which can be leveraged to create more accurate and personalized recommendations. This type of recommendation system considers not only the past behaviors and preferences of users but also the behaviors and preferences of their connections or friends in the network [8].
@@
Trust is a crucial concept in social networks and plays a vital role in recommendation systems [40, 3, 42, 61, 25, 33, 16]. The abstraction of social connections as ’trust’ is essential because it allows the system to consider implicit behaviors and preferences of users and their connections. For example, a user liking or sharing content, or having many mutual connections, can indicate a level of trust or shared interest, even if the users are not explicitly connected or have not directly interacted. This abstraction helps in capturing the nuances of social interactions and enables the recommendation system to provide more personalized and accurate suggestions. Additionally, the concept of trust incorporates a user’s social circle’s opinions and preferences, making recommendations more robust and comprehensive [48]. In this paper, we investigate trust connections, and the implications they have on a user’s preferences and the robustness to adversarial attacks.
@@
There are several types of recommendation systems, including content-based, collaborative filtering, hybrid, and others. Content-based recommendation systems suggest items based on the items’ characteristics and the users’ preferences [22]. On the other hand, collaborative filtering recommends items based on the preferences and behaviors of similar users [51]. Hybrid systems combine both approaches [2]. Trust-based collaborative filtering [10] is particularly interesting because it not only considers the preferences of similar users but also incorporates the trust relationships between users. This approach helps in mitigating some of the limitations of traditional collaborative filtering, such as the cold start problem and data sparsity, by leveraging the trust network to generate recommendations even when there is limited interaction data available [58]. We will explore this technique in this paper, but with a first principles approach to develop an understanding of how to utilize these different methods and information forms best.
@@
Outline. The paper is organized in the following manner. Section II gives an overview of prior work, and section III outlines our contribution. After that, in section IV we introduce the methodology and in section V we discuss the single information-based algorithms and subsequently multiple information-based algorithms in section VI. Validation of the findings on more datasets is given in section VII, comaprison against other state-of-the-art (SOTA) algorithms provided in section VIII and the impact of adversarial attacks is explored in section IX. A discussion of our main findings is given in section X.
@@ II.1 Traditional Approaches
Traditional trust-based recommendation systems mainly focus on incorporating explicit and implicit trust relationships into the recommendation process. For example, some approaches consider the ratings or preferences of a user’s trusted friends when generating recommendations. Messa et. al [38] delved into a trust-aware collaborative filtering approach. Their method distinctively pivots on explicit trust relationships. This means that the system directly harnesses expressed trust values (like friendship connections) to modulate the influence of a user’s peers on their recommendations.
@@ II.2 Cold-Start Approaches
Addressing the “cold start” problem remains a pivotal challenge for trust-based recommendation systems. The “cold start” issue surfaces when a new user joins the platform or a new item is added, and there’s minimal to no interaction data available for meaningful recommendation generation. In such scenarios, traditional collaborative filtering methods often fall short, as they rely predominantly on historical interaction data [39]. Various solutions have been proposed, including adaptive meta-learning [46], deep learning [56], and a demographic approach based on finding similarities between old and new users [41].
Another porpular approach is leveraging the trust network [30]. The assumption here is that even if a user hasn’t interacted with many items, their trusted connections might have. By leveraging this network, systems can extrapolate the new user’s potential interests based on their trusted peers’ preferences. One particularly noteworthy approach to leverage the trust network against the “cold start” problem was conceived by [27]. They introduced TrustWalker, a random walk model grounded in the trust network. It commences its journey from the target user and propagates through their trusted connections, and with some probability, swapping to other similar items that may have more information. In this way, TrustWalker harnesses both explicit and implicit trust connections along with skipping to different items, which lets “cold start” users with little information still receive “acceptable” recommendations.
@@ II.3 Machine Learning Approaches
Machine learning, over the past decade, has revolutionized numerous domains with its ability to extract patterns from vast data sets. Traditional methods of collaborative filtering were primarily linear. However, the dynamics of user preferences, trust relationships, and item attributes often exhibit nonlinear patterns that require more sophisticated models for accurate representation [37]. Hence, machine learning is a widely used tool in recommendation systems [29]. Bayesian [13, 15] and decision tree [49, 34] approaches are mostly used due to their simplicity [43]. Moreover, gradient-decent-based [9], matrix factorization-based [47, 55], non-linear collaborative filtering [26], and hidden Markov model-based [14] algorithms are also in use.
Neural collaborative filtering models underscore the potential of deep neural networks in handling the collaborative filtering challenge on implicit feedback data sets [24, 54]. They propose three different ML-based models for developing recommendation systems and show the models compete with other state-of-the-art approaches. Moreover, these approaches have also been used to tackle the cold-start problem [56, 53].
@@ II.4 Adversarial Defense Approaches
Low-rank defenses [12] against adversarial attacks are one type of adversarial defense strategy used in recommendation systems. Low-rank defenses entail transforming the user-item matrix into a low-rank matrix to decrease the impact of adversarial attacks. This method has been demonstrated to be effective in fighting against several sorts of adversarial attacks, including poisoning attacks and model extraction attacks. In recommendation systems, generative adversarial networks (GANs) can detect and fight against adversarial attacks [32]. GANs can be used to generate adversarial instances and teach the recommendation model to defend against them. This technique effectively defends against adversarial assaults such as poisoning and model extraction [11]. Moreover, methods based on these have shown great prospects in terms of modeling the long-tail distribution of recommendation systems [44]. Another study by [31] focused on how to accurately portray the representative rating for an item to mitigate the risk of an adversary adding fake ratings. The paper found that placing weights on network-related factors leads to increased robustness to adversaries.
@@ V.1.3 Monte-Carlo Random Walk
We now move to investigate recommendation systems that consider the entire trust graph. The Monte-Carlo (MC) Random Walk model works by starting at the node we are trying to produce a rating for and then performing a random walk. As the distance from the original node, $d$, increases, so does the probability that a random rating is returned. This random walk continues until a node with a rating is reached, or this random rating is produced. We run this experiment $k$ times, for some fixed $k$, for each node and take the mean rating from the nodes reached via random walk. Here $\alpha$ controls probability of making a random jump to a rated node instead of continuing the regular traversal. The pseudocode for this algorithm is given in Algorithm 1.
Algorithm 1 Monte Carlo Random Walk
1:function monteCarloRandomWalk($node$, $item$, $graph$, $ratings$, $k$)
2:     function traverse($node$, $distance=0$, $\alpha=0.05$)
3:         if $node\textbf{ in }ratings$ then
4:              return $ratings[node]$
5:         end if
[…]
7:              return $traverse(random.choice(ratings)])$
8:         end if
9:         $next\leftarrow graph[node]$
10:         if $|next.neighbours|==0$ then
11:              return None
12:         end if
13:         return $traverse(random.choice(next.neighbours)),d+1)$
14:     end function
15:     $results\leftarrow[]$
16:     for $i\textbf{ in }\text{range}(k)$ do
17:         $sampleRating\leftarrow traverse(node)$
18:         if $sampleRating==None$ then
19:              $results.push(random(1,5))$
20:         else
21:              $results.push(sampleRating)$
22:         end if
23:     end for
24:     return $mean(ratings)$
25:end function
@@ V.1.4 Jaccard Majority of Majorities (JMoM) and Weighted Average (Jaccard WA)
These recommenders work by considering every other node, which has rated item i, in the graph. We then perform a summation of the Jaccard values for every node corresponding to each rating value. In the Majority of Majorities version, we then take the rating that has received the highest contribution of Jaccard scores. However, in the Weighted Average version, we instead take the weighted average of these ratings, weighted by the contribution of Jaccard scores. We will call these techniques JMoM and WA moving forward. (These methods are inspired by the popular majority based opinion diffusion models, cf. [18, 5, 59].). The pseudocode for weighted average is given in Algorithm 2.
Algorithm 2 Weighted-Average
1:function weightedAverage($node$, $item$, $graph$, $ratings$, $similarityFunction$)
2:     $buckets\leftarrow Array(5)$
3:     for $otherVertex\textbf{ in }graph$ do
4:         if $otherVertex\textbf{ in }ratings\textbf{ \&\& }otherVertex\neq node$ then
5:              $idx\leftarrow ratings[otherVertex][item]-1$
6:              $buckets[idx]\leftarrow buckets[idx]+similarityFunction(current,otherVertex)$
7:         end if
8:     end for
9:     if $potentialRatings.sum==0$ then
10:         return $ratings$ randomInteger(1, 5)
11:     end if
12:     $WA\leftarrow 0$
13:     for $i\textbf{ in }range(5)$ do
14:         $WA\leftarrow WA+buckets[i]*(i+1)$
15:     end for
16:     return $WA$
17:end function
@@ VII Validation With Other Data Sets
Jaccard Item-Jaccard WA 1.00 0.26 0.66 0.08 0.56 0.28
@@
It appears the Jaccard Item-Jaccard WA recommendation algorithm performed best across all data sets with the lowest MAE value. The algorithm consistently outperformed other algorithms in each of the data sets - Epinions, FilmTrust, and CiaoDVD - with MAEs of 1.00, 0.66, and 0.56 respectively. This algorithm’s Mean MAE ($\mu_{MAE}$) was significantly lower than other algorithms.
@@
VIII Comparison Against SOTA Algorithms
Here, we compare our results with other previously established methods. We compared against algorithms like SoRec[35], SoReg [36], TrustSVD[20], TrustMF [57], and APURTJ [1]. APURTJ stands for Adjusted Triangle, User Rating Preferences behavior and Jaccard similarity. This algorithm outperforms the other methods.
The optimal experimental settings for each method are suggested by previous works [52]. The settings are: (1) SoRec: the number of latent features $d=5$ for SoRec and $\lambda_{c}=1.0,0.01,1.0$ corresponding to FilmTrust, CiaoDVD and Epinions respectively; (2) SoReg: the number of latent features d = 5 for SoReg and $\beta=0.1$ for the all; (3) TrustMF: : the number of latent features $d=5$ and $\lambda_{t}=1$ ; (4) TrustSVD: the number of latent features $d=5$ for TrustSVD and $\lambda=0.1$, $\lambda_{t}=0.9$ for FilmTrust, $\lambda=0.6$, $\lambda_{t}=0.5$ for Epinions, and $\lambda=0.5$, $\lambda_{t}=1.0$ for CiaoDVD.
Based on the results given in Table 12, we observe that while our algorithms do not necessarily outperform all other methods, their performance is satisfactory, especially considering their simplicity and scalability. It is worth to emphasize that the main goal of this paper is a systematic study of the impact of various sources of information on the recommendation quality, rather than providing a better recommendation algorithm. However, the above comparison has been conducted to complement the understanding of the reader.
Table 12: Mean of MAE for SOTA Algorithms on Different Data Sets
Algorithm Epinions FilmTrust CiaoDVD
Jaccard Item-Jaccard WA 1.00 0.66 0.56
SoReg 0.96 0.64 0.9
SoRec 0.92 0.72 0.81
TrustSVD 0.82 0.61 0.72
TrustMF 0.83 0.63 0.76
ATURPJ - 0.5 0.55
Universal Random 1.30 0.89 0.72
Random 1.67 1.34 1.82
IX Robustness Against Adversarial Attacks
@@ X Discussion
It is very important to note that the goal of our study is not focused on surpassing state-of-the-art techniques; our study is intended to be a methodical investigation of the function of different factors in recommendation systems. Numerous sophisticated methods, like deep learning-based strategies, make use of complex architectures that maximize efficiency but frequently mask the distinct contributions of various elements that can be extracted from the social networks. We allow for a controlled study of the impact of incremental information on recommendation quality by keeping the setting simple. This allows us to isolate and better understand the contribution of each parameter, which would be challenging in highly complex setups. Interestingly, our model also performs comparatively well against other methods, showing the power of graph structure information that are often ignored, while being very simple and scalable.
The computational complexity of various similarity metrics and recommendation methods is analyzed in terms of the number of users ($N$), the number of neighbors for a given user ($n$), the number of items reviewed by a user ($i$), and the complexity of a similarity metric ($s$). The complexity of computing a metric for one pair of items or users varies: the Jaccard Index has a complexity of $O(n)$, the Item-Jaccard Index and Item Rating Difference both have a complexity of $O(i)$, while the Intra-Item Jaccard requires $O(N)$ computations. When computing recommendations for a single user, the Weighted Average approach operates at $O(N*s)$, the Mean/Median of Neighbors method has a complexity of $O(n)$, and the Random Walk Monte-Carlo method scales as $O(\frac{1}{\alpha})$. These methods are much faster than sophisticated collaborative filtering methods like LightGCN [23] or Neural Collaborative Filtering [24].
@@
Paras Stefanopoulos: Conceptualization, Formal Analysis, Methodology, Software, Validation, Writing - Original draft preparation. Sourin Chatterjee: Software, Validation, Writing - Original draft preparation. Ahad N. Zehmakan: Conceptualization, Methodology, Supervision, Validation, Writing - Review.